what will be the molar concentration of a 500 mL solution that has 22.5g of hydrogen iodide dissolved in it?
1
Expert's answer
2020-06-04T10:37:52-0400


Hydrogen iodide HI


"n=frac{m}{M}=\frac{22.5}{127.913}=0.1759"


"C=\frac{n}{V}=\frac{0.1759}{0.5}=0.35"